OpenStudy (anonymous):

HELP me with my calculus assignmat. Please help me and get a medal. Just the first ten question. 1 to 10.

OpenStudy (anonymous):

Show your work.

OpenStudy (imstuck):

the first one looks like this when you rewrite without the radicals:\[\frac{ x ^{\frac{ 9 }{ 7 }} }{ x ^{\frac{ 6 }{ 5 }} }\]In order to combine those you need a common denominator between the 7 in the numerator's power and the 5 in the denominator's power. That common denominator is 35. So like this:

OpenStudy (imstuck):

\[\frac{ x ^{\frac{ 45 }{ 35 }} }{ x ^{\frac{ 42 }{ 35 }} }\]the rule then is to subtract the power in the denominator from the power in the numerator, like this:

OpenStudy (imstuck):

\[x ^{\frac{ 45 }{ 35 }-\frac{ 42 }{ 35 }}\]which gives you \[x ^{\frac{ 3 }{ 35 }}\]which as a single radical is this:\[\sqrt[35]{x ^{3}}\]

OpenStudy (imstuck):

the second one you will begin by factoring out a single x, leaving you with\[x(6x ^{2}-17x+5)\]Factoring the remaining polynomial gives you factors of\[(x-\frac{ 5 }{ 2 })(x-\frac{ 1 }{ 3 })\]so factoring completely you get this\[x(x-\frac{ 5 }{ 2 })(x-\frac{ 1 }{3 })\]

OpenStudy (imstuck):

Can't find the zeros of #3 because they are nonreal complex. I cannot even seem to find them on my graphing calculator. It seems to me that the range though goes from negative infinity to 49.12